OpenStudy (anonymous):

write the related series for the finite sequence 3,7,11...........,27. then find the sum

OpenStudy (anonymous):

S(n) = 3+4(n-1) = \[\sum_{0}^{7}(3+4(n-1))\] Sum = 30*3 + 15 = 105

OpenStudy (anonymous):

3,7,11,15,19,27

OpenStudy (anonymous):

3,7,11,15,19,23,27=105